OpenStudy (jacky90):

Working together, it takes two computers 10 minutes to send out a company's email. If it takes the slower computer 35 minutes to do the job on its own, how long will it take the faster computer to do the job on its own? Do not do any rounding. @sshayer

OpenStudy (phi):

use the idea rate * time = 1 job and we need the rate = jobs per minute in this problem, one of the computers does one job in 35 minutes that means its rate is 1/35 call the other rate x (for unknown) then together (we add their rates) we can say (1/35 + x) * 10 = 1 can you solve for x?

OpenStudy (sshayer):

\[\frac{ 1 }{ x }=\frac{ 1 }{ 10 }-\frac{ 1 }{ 35 }=\frac{ 7-2 }{ 70 }=\frac{ 5 }{ 70 }=\frac{ 1 }{ 14 }\] x=?